OpenStudy (anonymous):

[6.02] For the following system, if you isolated y in the first equation to use the Substitution Method, what expression would you substitute into the second equation? 4x − y = 8 2x + 3y = 9 A: 8 − 4x B: 8 + 4x C: −8 + 4x D: −8 − 4x

OpenStudy (anonymous):

Can you tell me what you will get from 4x − y = 8 if you decided to isolate y in this equation?

OpenStudy (anonymous):

Love the ponies <3 And I honestly forget how to do these. My first reaction is too minus the y but that doesn't really do anything. How do i isolate?

OpenStudy (anonymous):

isolating y mean that we want y by itself either on the left side or the right side. Sine y is on the left side we want to keep it there so we dont have to do more work. That doesn mean we cant have y on the right side it's just easier to keep it on the left side. Noe we want to get rid of everything else on the left side except for y. That means we want to get rid of 4x and in order to do that we want to subtract 4x from it. But remember what we do on one side we have to do it to the other side as well so we have to subtract 4x to both sides. This is how it looks like 4x-4x − y = 8-4x the two 4x cancel on the left side and you are left with − y = 8-4x

OpenStudy (anonymous):

Do you follow along?